OpenStudy (anonymous):

Solve the system using matrices. {x+2y=-1 {-x-3y=-2 Can someone please help? I don't get this at all.

hero (hero):

http://youtu.be/YnjqgjefRyg?t=4m42s

hero (hero):

Ax = B A = [1 2] [-1 -3] x = B/A x = 1/A *B x = A^(-1)*B Find the inverse of A then set x = A^(-1)*[-1] [-2]

hero (hero):

@zozeebo, have you found the inverse of A yet?

OpenStudy (anonymous):

I did, and I figured out the answer! Thank you!

hero (hero):

Congrats